Comparative financial statements for Weaver Company follow:

Weaver Company
Comparative Balance Sheet
at December 31
This Year Last Year
Assets
Cash and cash equivalents $ 10 $ 20
Accounts receivable 590 370
Inventory 170 235
Prepaid expenses 9 7
Total current assets 779 632
Property, plant, and equipment 680 570
Less accumulated depreciation 75 60
Net property, plant, and equipment 605 510
Long-term investments 2 47
Total assets $ 1,386 $ 1,189
Liabilities and Stockholders' Equity
Accounts payable $ 390 $ 285
Accrued liabilities 35 50
Income taxes payable 84 77
Total current liabilities 509 412
Bonds payable 370 270
Total liabilities 879 682
Common stock 346 450
Retained earnings 161 57
Total stockholders’ equity 507 507
Total liabilities and stockholders' equity $ 1,386 $ 1,189
Weaver Company
Income Statement
For This Year Ended December 31
Sales $ 870
Cost of goods sold 485
Gross margin 385
Selling and administrative expenses 207
Net operating income 178
Nonoperating items:
Gain on sale of investments $ 3
Loss on sale of equipment (1 ) 2
Income before taxes 180
Income taxes 54
Net income $ 126

During this year, Weaver sold some equipment for $18 that had cost $48 and on which there was accumulated depreciation of $29. In addition, the company sold long-term investments for $48 that had cost $45 when purchased several years ago. Weaver paid a cash dividend this year and the company repurchased $104 of its own stock. This year Weaver did not retire any bonds.

Required:

1. Using the direct method, adjust the company’s income statement for this year to a cash basis.

2. Using the information obtained in (1) above, along with an analysis of the remaining balance sheet accounts, prepare a statement of cash flows for this year.

Complete this question by entering your answers in the tabs below.

  • Required 1
  • Required 2

Using the direct method, adjust the company’s income statement for this year to a cash basis. (Adjustment amounts that are to be deducted should be indicated with a minus sign.)

Weaver Company
Direct Method of Determining the Net Cash Flows from Operating Activities
Adjustments to a cash basis:
Adjustments to a cash basis:
Selling and administrative expenses
Adjustments to a cash basis:
Income taxes
Adjustments to a cash basis:
     
  • Required 1
  • Required 2
  • Complete this question by entering your answers in the tabs below.

  • Required 1
  • Required 2
  • Using the information obtained in (1) above, along with an analysis of the remaining balance sheet accounts, prepare a statement of cash flows for this year. (Cash outflows and amounts to be deducted should be indicated with a minus sign.)

    Weaver Company
    Statement of Cash Flows
    For This Year Ended December 31
    Operating activities:
    Cash received from customers
    Less cash disbursements for:
    Total cash disbursements
    Investing activities:
    Financing activities:
    Beginning cash and cash equivalents
    Ending cash and cash equivalents   
  • Required 1
  • Required 2
Answer #1
Cash Flow Statement
Direct Method
Cash Flow from Operating Activities
Cash Collected from Customers $             650.00 =870+370-590
Cash paid for inventory $            -315.00 =-(485+285-390+170-235)
Cash paid for Selling and administrative expenses $            -180.00 =-(207+50-35+9-7)+44
Cash paid for Income tax $              -47.00 =-(54+77-84)
Cash from Operating Activities $             108.00
Cash Flow Statement
Indirect Method
Cash Flow from Operating Activities
Net Income $             126.00
Adjustments
Depreciation $           44.00 =75-(60-29)
Gain on sale of Investment $            -3.00
Loss on sale of Equipment $              1.00
Changes in Current Assets/ Current Liabilities
Increase in Accounts Receivable $       -220.00 =370-590
Decrease in Inventory $           65.00 =235-170
Inrease in prepaid expenses $            -2.00 =7-9
Increase in Accounts Payable $         105.00 =390-285
Decrease in Accrued Liabilities $          -15.00 =35-50
Increase in Income Tax Payable $              7.00 =84-77
Total Adjustments $              -18.00
Cash from Operating Activities $             108.00
Cash flow from Investing Activities
Sale of Investments $                48.00
Sale of Equipment $                18.00
Purchase of Equipment $            -158.00 =570-48-680
Net cash used In investing activities $             -92.00
Cash flow from Financing Activities
Repurchase of Common Stock $            -104.00 =346-450
Issue of Bonds $             100.00 =370-270
Dividend Paid $              -22.00 =-(57+126-161)
Net cash used in financing activities $             -26.00
Increase in Cash $             -10.00
Opening Balance of Cash $               20.00
Closing Balance of Cash $               10.00